OpenStudy (anonymous):

Complex Analysis: Residue theory Verify the integral formula with aid of residue: integral of (((x^2)+1)/((x^4)+1)) dx from 0 to infinity = PI / (Square-root of 2). My thinking: Let: a1 = square-root of i a2 = square-root of -i I can see the solution: (1/2) (2PI * i) * res (a1) + res (a2) Where: (1/2) is half of circle (2PI*i) is the circle stuff Problem: I can not seem to crack the res(a1) and res(a2). Where are you KingGeorge? THanks. Regards DL

OpenStudy (anonymous):

\[ a_1=\sqrt{i}=\frac{1+i}{\sqrt{2}} \]

OpenStudy (anonymous):

Now find \[ \lim{x->a1} \left(\frac{\left(x-\frac{1+i}{\sqrt{2}}\right) \left(x^2+1\right)}{x^4+1}\right)= -\frac{i}{2 \sqrt{2}} \] Do the steps necessary to prove that

OpenStudy (anonymous):

Thanks so much EliasSaab for your help and time. It is appreciated. As I am self learning, it is a little bit of difficulty to follow: sqrt(i) = (1+i)/sqrt(2). Also, please let me know if I am err on the res(a2), also, the (1/2) part as well. Basically, is my thinking generally correct or not? In the meantime, I will try your suggestions. Again, thanks so much for your help and time. It is wonderful. Just another question, do you know of anything about k-factor things since I am very interested in this k-factor (loading analysis); but I could not find anything about it at all. Thanks again;

OpenStudy (anonymous):

Thanks EliasSaab, I can work out the srqt(i) you gave me. However, I wonder how in the world do you cam up with that identity. Thanks.

OpenStudy (anonymous):

Hi EliasSaab, Thanks for your time. I worked out the limit: ((x^2)+1)/(x+sqrt(i)) * (x^2+i). as x->sqrt(i) = (1+i)/2. (1/2)(2PI*i)*(-i/2*srqt(2)) = - (PI* (i^2))/sqrt(2) = PI/sqrt(2). So things are fine now. However, I need to know if my thinking was err since I did not do res (a2). Please let me know your comments, ideas or questions. Thanks so much for your helps and time. It is appreciated. Regards DL

OpenStudy (anonymous):

Write \[ i = e^{ \frac {i \pi} 2 }\\ \sqrt i= \left( e^{ \frac {i \pi} 2 }\right)^{\frac 1 2}= e^{ i\frac \pi 4} =\cos(\pi/4) + i \sin(\pi/4)= \frac 1 {\sqrt 2} + i \frac 1 {\sqrt 2} \]

OpenStudy (anonymous):

Thanks a lot. eliassaab. Simply amazing @ the sqrt(i)=(1+i)/sqrt(2).